Apply definition of antiderivative and find area under the curve of f(x) = x^1/2 between x=0 and

x=1


1
Expert's answer
2022-03-10T18:16:52-0500

Solution


The required area is calculated as:


"A = \int\limits_a^b {f\left( x \right)\,} dx\"


Here "f(x)=\sqrt{x}" and the limits are "a=0" and "b=1"


Therefore, the required area is


"A = \int\limits_0^1 {\sqrt x \,} dx\"


"A=[\frac{x^\frac{3}{2})}{\frac{3}{2}}]_{0}^{1}"


"A = \frac{2}{3}\left[ {{x^{{\textstyle{2 \over 3}}}}} \right]_0^1\"


"A = \frac{2}{3}\left[ {{{\left( 1 \right)}^{{\textstyle{2 \over 3}}}} - {{\left( 0 \right)}^{{\textstyle{2 \over 3}}}}} \right]\"


"A = \frac{2}{3}\left[ {1 - 0} \right]\"


"A = \frac{2}{3}" square units


The shaded area below shows the required area.